A cylindrical drum has a radius of 0.45 m and is initially at rest. It is then given an angular acceleration of 0.40 rad s−2. At time t = 8.0 s calculate (i) the angular speed of the drum, (ii) the centripetal acceleration of a point on the rim of the drum, (iii) the tangential acceleration at that point, and (iv) the resultant acceleration at that point.
1
Expert's answer
2019-04-01T10:29:57-0400

(i)


"\omega = \alpha t=(0.4)(8)=3.2 rad/s"

(ii)


"a_c=\omega^2r=(3.2)^2 0.45=4.6 m/s^2"

(iii)


"a_t=\alpha r=(0.4)(0.45)=0.18 m/s^2"

(iv)


"a= \sqrt{0.18^2+4.6^2} =4.6 m/s^2"